OpenStudy (anonymous):

Toss 3 fair coins and let x equal the number of heads observed. a. Identify the sample points associated with this experiment and assign a value of x to each sample point b. Calculate p(x) for each value of x c. Construct a probability histogram for p(x) d. What is P(x = 2 or x = 3)?

OpenStudy (anonymous):

Well, \(X\) can be from \(0\) to \(3\).

OpenStudy (anonymous):

There are \(2^3=8\) possible outcomes. There is \(1\) outcome for \(X=0\), \(3\) for \(X=1\), \(3\) for \(X=2\), and \(1\) for \(X=3\).

OpenStudy (anonymous):

Hi Wio, thanks for helping. I'm not sure how to approach this problem. I understand the 2^2 = 8 part. I think we need to make a tree for this?

OpenStudy (anonymous):

By tree, do you mean pascal's triangle?

OpenStudy (anonymous):

Actually, do we just list it out like this? HHH HTH HTT HHT THH TTH THT TTT those are the sample points, correct?

OpenStudy (anonymous):

Yeah, sure

OpenStudy (anonymous):

Ok, so each of those are X?

OpenStudy (anonymous):

P(1) = 3/8 p(2) = 4/8 p(3) = 1/8 am I doing this correctly?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

you need \(p(0)\)

OpenStudy (anonymous):

p(0) = 1/8 ?

OpenStudy (anonymous):

yes, and your \(p(2)\) is wrong

OpenStudy (anonymous):

p(2) = 3/8 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you, i think i can solve the rest